Over a 44-year period an asset had an arithmetic return of 11.8 percent and a geometric return of 9.7 percent. Using Blume’s formula, what is your best estimate of the future annual returns over 7 years? 12 years? 22 years? (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.) Future annual returns 7 years % 12 years % 22 years %
Answer #1

Using Blume's formula Return at time(T) =(T-1)/(N-1)*Geometric Return+(N-T)/(N-1)*Average Return

Return in 7 years=(7-1)/(44-1)*9.7%+(44-7)/(44-1)*11.80% = 11.51%
Return in 12 years =(12-1)/(44-1)*9.7%+(44-12)/(44-1)*11.80%=11.26%
Return in 22 years =(22-1)/(44-1)*9.7%+(44-22)/(44-1)*11.80% =10.77%